Security ABC has a price of $35 and a beta of 1.5. The risk-free rate is 5% and the market risk premium is 6%.
a)Explain the terms beta and market risk premium.
b)What is the market portfolio?
c)According to the CAPM, what return do investors expect on the security?
d)Investors expect the security not to pay any dividend next year.
e)At what price do investors expect the security to trade next year?
f)At what price do investors expect the security to trade next year, if the expected dividend next year is $2 instead of zero?
Under the CAPM, expected returns depend only on the exposure to systematic (market) risk. Illustrate the concept of systematic risk through examples and explain the logic of this result.
Answer #1

A) Beta is a measure of systematic risk which measures the risk in relative terms. For example if any security beta is 1.2 and market beta is 1, then if market moves up by 10% the upward movement in the security would 10*1.2= 12%, similarly if the market went down by 10 percent then the security will go down by 12%. Beta measures the relative risk of any security.

Market Risk premium : Market risk premium is the difference between the return from the market portfolio and risk free rate. It is the excess premium required to invest in risky projects. Market risk premium is also equal to the slope of the security market Line.

B) Market portfolio : Creating a market portfolio for a small Invetsor might not be efficient cost wise but large investors can do that. Market portfolio is a portfolio in which all securities are present in same weight as they are in the market. For example if you want to replicate the performance of S&P 500, you have to buy all the constituents in the same weight as they are in S&P 500.

C).

According to CAPM model required rate of return is =

= risk free rate + Beta *(Market risk premium)

=5+ 1.5*6 = 14%

D)

if the security does not pay any dividend then the next year price according to CAPM model should be

35*(1.14) =$39.90